# Training Large Neural Networks With Low-Dimensional Error Feedback

### v3

@@ Abstract
Training deep neural networks typically relies on backpropagating high-dimensional error signals—a computationally intensive process with little evidence supporting its implementation in the brain. However, since most tasks involve low-dimensional outputs, we propose that low-dimensional error signals may suffice for effective learning. To test this hypothesis, we introduce a novel local learning rule based on Feedback Alignment that leverages indirect, low-dimensional error feedback to train large networks. Our method decouples the backward pass from the forward pass, enabling precise control over error signal dimensionality while maintaining high-dimensional representations. We begin with a detailed theoretical derivation for linear networks, which forms the foundation of our learning framework, and extend our approach to nonlinear, convolutional, and transformer architectures. Remarkably, we demonstrate that even minimal error dimensionality—on the order of the task dimensionality—can achieve performance matching that of traditional backpropagation. Furthermore, our rule enables efficient training of convolutional networks, which have previously been resistant to Feedback Alignment methods, with minimal error. This breakthrough not only paves the way toward more biologically accurate models of learning but also challenges the conventional reliance on high-dimensional gradient signals in neural network training. Our findings suggest that low-dimensional error signals can be as effective as high-dimensional ones, prompting a reevaluation of gradient-based learning in high-dimensional systems. Ultimately, our work offers a fresh perspective on neural network optimization and contributes to understanding learning mechanisms in both artificial and biological systems.
@@
Conventional wisdom holds that training large neural networks requires propagating precise, high-dimensional error gradients through every neuron—a process lacking biological plausibility. Our work demonstrates that effective learning can instead occur with low-dimensional error signals, whose dimensionality scales with task complexity rather than network size. By developing a novel learning paradigm that employs local plasticity rules to adaptively align feedback signals with task-relevant errors, we show theoretically and empirically that even minimal error dimensionality achieves competitive performance in complex neural architectures. These findings challenge current assumptions about neural optimization and highlight fundamental constraints that can govern learning processes in both artificial and biological neural systems.
@@ 1.   Introduction
The brain, despite its vast complexity, operates within the constraints of a low-dimensional world. Every movement—lifting a hand, turning a head—follows finite joint constraints. Navigation unfolds in modest spatial dimensions, and object recognition is drawn from a limited set of familiar categories. Even when possibilities seem vast, they are dwarfed by the sheer number of neurons and synapses in any brain region. Artificial neural networks reflect this paradox: the number of neurons in each layer often far exceeds the task-relevant variables. This contrast between task and internal representation dimensionality challenges us to rethink how we train such large networks and explore new learning paradigms.
Modern deep learning methods rely on high-dimensional gradients that propagate backward through layers, ensuring a precise credit assignment to each parameter [1, 2]. However, this process is computationally intensive and biologically implausible, given the lack of direct evidence for backpropagation in the brain [3, 4]. If the loss gradient does not propagate back through the network, it must be obtained through alternative routes [5]. In this case, there is no reason to assume that each neuron receives detailed information about the local gradient. Moreover, the brain likely operates with low-dimensional error signals, reflecting the inherently low-dimensional nature of tasks. This raises an essential question: Can large neural networks be trained efficiently with partial information about the loss gradient carried by low-dimensional error signals? If error dimensionality is tied to task complexity rather than network size, it could revolutionize how we train large networks and improve our understanding of biological learning.
Despite the importance of this question, the effect of low-dimensional error feedback on learning and neural representations was never properly explored. The strong coupling of feedforward and backward passes in backpropagation, where both processes use the same synapses, has limited investigations into alternative training methods. Some approaches, such as Feedback Alignment [5], were proposed to address the problem of weight transposition in backporpagation, the necessity of full gradients for effective learning has not been questioned.
In this study, we propose that high-dimensional error propagation is not strictly necessary for training deep, overparameterized networks. We introduce a novel learning rule based on Feedback Alignment that decouples forward and backward passes, enabling precise control over the dimensionality of error signals. Our method exploits the low-dimensional nature of real-world tasks to train large networks efficiently without compromising their representational capacity, achieving performance on par with backpropagation. A theoretical analysis in linear settings highlights the potential pitfalls of naive low-dimensional error propagation and shows that a simple local online learning rule can recover backpropagation-level performance. Extending this approach to nonlinear, convolutional, and transformer-based networks, we demonstrate that learning with low-dimensional error signals is feasible and universal. Finally, we explore how error dimensionality influences neural representations by examining its effects on receptive fields in an early visual system model, underscoring the broader implications of our findings.
@@ 2.   Background and related work
We consider a multilayered perceptron with $L$ layers, each layer $l$ computes its output as $\boldsymbol{h}_{l}=f(W_{l}\boldsymbol{h}_{l-1})$, where $W_{l}$ is the weight matrix and $f$ is an element-wise activation function. The input to the network is $\boldsymbol{h}_{0}=\boldsymbol{x}$, and the final network output is $\hat{\boldsymbol{y}}=f_{L}(W_{L}\boldsymbol{h}_{L-1})$, which approximates the target $\boldsymbol{y}$. The task dimensionality, denoted $d$, is at most the number of components in $\boldsymbol{y}$ and $\hat{\boldsymbol{y}}$. Training the network involves minimizing a loss function $\mathcal{L}(\boldsymbol{y},\hat{\boldsymbol{y}})$ by adjusting the weights $\{W_{l}\}$. The error signal at the output layer, $\delta_{L}=\frac{\partial\mathcal{L}}{\partial\hat{\boldsymbol{y}}}$, is a $d$-dimensional vector, typically much smaller than the number of neurons in the hidden layers.
Backpropagation (BP) [1] is the standard approach for training neural networks. It propagates the error backward through the network using $\delta_{l}=W_{l+1}^{T}\delta_{l+1}\odot f^{\prime}(W_{l}\boldsymbol{h}_{l-1})$, and updates the weights using $\Delta W_{l}=-\eta\delta_{l}\boldsymbol{h}_{l-1}^{T}$, where $\eta$ is the learning rate. However, this method requires an exact transposition of the forward weights, $W_{l+1}^{T}$, which is biologically implausible [6, 7]. Moreover, BP tightly couples the error propagation with the forward pass, limiting the ability to explore how different properties of the error signal affect learning dynamics.
Feedback Alignment (FA) [5] was proposed to address the biological limitations of BP by replacing $W_{l+1}^{T}$ with a fixed random matrix $B_{l}$. The error is computed as:
@@
$\Delta B_{l}=-\eta\boldsymbol{h}_{l-1}\delta_{l}^{T}-\lambda B_{l},\quad\Delta W_{l}=-\eta\delta_{l}\boldsymbol{h}_{l-1}^{T}-\lambda W_{l},$ (2)
where $\lambda$ is a regularization parameter. Although this adaptive approach improves performance by better aligning forward and backward weights, it still requires high-dimensional error signals and struggles to match BP performance in complex architectures like CNNs. Furthermore, in Section 3, we show that this approach fails when the matrix $B$ is low-rank and the dimensionality of the error is constrained.
@@
In this work, our objective is to systematically investigate how constraining the dimensionality of the error signal affects the training and performance of large neural networks. To this end, we introduce a novel learning scheme that enables deep layers to be learned using low-rank feedback matrices, which allow flexible control over the dimensionality of the loss gradients in each layer (Figure 1).
Our main contributions are:
  1. 1.
    We present a novel local learning rule for low-rank feedback weights, which matches BP performance while requiring minimal error signals. We provide a detailed derivation of the learning dynamics in a simple linear case, establishing the foundation for why and how the feedback weights should be learned.
  2. 2.
    We demonstrate that large and deep nonlinear networks can efficiently learn nontrivial datasets using low-dimensional error signals and highlight that the feedback dimensionality should scale with the task rather than the size of the network.
  3. 3.
    We successfully train complex yet highly useful architectures, such as convolutional networks and visual transformers, using low-dimensional feedback.
  4. 4.
    We study a simplified model of the ventral visual system and show that the emerging receptive fields depend on the error-feedback dimensionality. These findings offer new insights into the relationship between learning mechanisms and biological neural representations.
In the final section, we discuss the broader implications of our results for both neuroscience and machine learning.
Figure 1: Illustration of different approaches for propagating error to hidden layers. From left to right: Backpropagation (BP) propagates error using the exact transpose of the forward weights. Feedback Alignment (FA) replaces the transposed weights with fixed random feedback matrices, which gradually align with the forward weights during training. Our work framework provides low-dimensional feedback (LDF) using low-rank matrices, allowing full control of the dimnsionality of the gradients in each layer. Finally, direct low-dimensional feedback (dLDF) extends our LDF architecture by enabling error signals to bypass intermediate layers, propagating directly from the output layer or other non-adjacent layers.
@@ 3.   Low-dimensional error feedback in linear networks
We consider a simple linear transformation problem with a low-dimensional structure, $\boldsymbol{y}=A\boldsymbol{x}$. Here, $\boldsymbol{x}\in\mathbb{R}^{n}$ represents the $n$-dimensional input, and $\boldsymbol{y}\in\mathbb{R}^{m}$ represents the target. The matrix $A$ is a rank-$d$ matrix defined as $A=\sum_{j=1}^{d}\boldsymbol{u}_{j}\boldsymbol{v}_{j}^{T}$, where $\boldsymbol{u}_{j}\in\mathbb{R}^{n}$ and $\boldsymbol{v}_{j}\in\mathbb{R}^{m}$ are random Gaussian vectors, and we assume $d\ll n$. Our data set consists of $p$ training samples $\{\boldsymbol{x}^{\mu},\boldsymbol{y}^{\mu}\}_{\mu=1}^{p}$, with each input vector $\boldsymbol{x}^{\mu}$ being i.i.d. according to the standard normal distribution $\boldsymbol{x}^{\mu}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{1})$. The labels are given by $\boldsymbol{y}^{\mu}=A\boldsymbol{x}^{\mu}+\boldsymbol{\xi}^{\mu}$, where $\boldsymbol{\xi}^{\mu}$ is additive Gaussian noise with zero mean and unit variance.
@@
We apply Feedback Alignment (FA) to update $W_{1}$, which does not have direct access to the loss gradient. Instead of backpropagating the error through $W_{2}^{T}$, we use a fixed low-rank feedback matrix $B$. This provides an alternative pathway for propagating the error signal to $W_{1}$.
@@
$\Delta W_{1}^{\mu}$ $=\eta B(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu})\boldsymbol{x}^{\mu T},\text{ and }$ (4)
$\Delta W_{2}^{\mu}$ $=\eta(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu})\boldsymbol{x}^{\mu T}W_{1}^{T}.$
@@ 3.1 Learning dynamics
Our analysis extends the framework established by Saxe et. al. [16] to incorporate indirect feedback with constrained dimensionality. We begin by characterizing the task across the $p$ data points through the input-output covariance matrix, $\Sigma_{io}=\frac{1}{p}\sum_{\mu=1}^{p}\boldsymbol{y}^{\mu}(\boldsymbol{x}^{\mu})^{T}$, which captures the correlation between input vectors $\boldsymbol{x}$ and output vectors $\boldsymbol{y}$. Performing Singular Value Decomposition (SVD), we obtain $\Sigma_{io}=USV^{T}$, where $U\in\mathbb{R}^{m\times m}$ and $V\in\mathbb{R}^{n\times n}$ contain the left and right singular vectors, respectively, and $S\in\mathbb{R}^{m\times n}$ is a rectangular diagonal matrix of singular values. For sufficiently large $p$, the first $d$ singular values in $S$ correspond to the prominent directions in the data (i.e., the singular values of $A$), while the remaining singular values are $O(1/\sqrt{p})$ and dominated by noise.
@@
$\tau\frac{d\bar{W}_{1}}{dt}=\bar{B}(S-\bar{W}_{2}\bar{W}_{1}),\quad\tau\frac{d\bar{W}_{2}}{dt}=(S-\bar{W}_{2}\bar{W}_{1})\bar{W}_{1}^{T},$ (5)
@@
where $\eta=dt/\tau$. This continuous form captures the dynamics of the learning process, allowing us to study it from a dynamical systems perspective. By analyzing these equations, we can identify fixed points and evaluate their stability, providing insight into how the network converges and learns under constrained feedback. Figure 2, shows how the singular vectors of $W_{2}W_{1}$ align to the corresponding singular vectors of $\Sigma_{io}$
@@
Figure 2: Learning dynamics and component alignment in linear networks. (a) Schematic of the network architecture with input dimension $n=128$, hidden layer size $k=64$, and output dimension $m=64$. The feedback matrix $B$ is factorized as $B=QP$ and constrained to rank $r$. (b, c) Comparison of theoretical predictions (dashed) and numerical simulations (solid) for low-rank Feedback Alignment (FA), training only $Q$[10], with $r=m=64$ and $r=8<m$, respectively. The $y$-axis shows the alignment of singular vectors between $W_{2}W_{1}$ and $\Sigma_{io}$. (d, e) Same as (b, c) but training both and $Q$ and $P$ using Eqs (7) and (8) . (f) When $P$ is not trained, the singular modes align on average (bold), but the top $r$ components fail to recover fully. (g) Training $P$ to align with the error ensures full recovery of the top $r$ singular components, improving convergence and alignment.
Stationary solutions for the training.
@@
Training halts when the right-hand side of the weight update equations (5) vanishes, indicating that the dynamics have reached a stable fixed point. At this fixed point, the update equation for $\bar{W}_{1}$ leads to the condition:
$\bar{B}(S-\bar{W}_{2}\bar{W}_{1})=0\implies S_{jj}\boldsymbol{B}_{:,j}=\sum_{i=1}^{m}\boldsymbol{B}_{:,i}(\bar{W}_{2}\bar{W}_{1})_{ij}\quad\forall j,$ (6)
where $\boldsymbol{B}_{:,j}\in\mathbb{R}^{k}$ is the $j$-th column of $\bar{B}$, and $S_{jj}$ is the $j$-th singular value of $\Sigma_{io}$. This equation indicates that, at the stationary point, the weight products $\bar{W}_{2}\bar{W}_{1}$ must align with the singular modes of the data.
However, since $B$ is of rank $r$, the feedback matrix $\bar{B}$ can only span at most $r$ independent directions. If $r=m$, the system has enough feedback dimensionality to align perfectly with the singular values in $S$, recovering the full structure of $\Sigma_{io}$ as demonstrated in Saxe et. al. [16]. In this case, the training successfully converges to a unique solution where $\bar{W}_{2}\bar{W}_{1}=S$ (Figure 2b).
Crucially, when $r<m$, the feedback matrix $\bar{B}$ lacks the sufficient rank to fully capture the $m$ independent directions in $S$. As a result, eq. (6) becomes under-determined, leading to potentially infinite solutions. This means that the trained weights $\bar{W}_{2}\bar{W}_{1}$ may not align with the true data structure encoded in $\Sigma_{io}$ (Figure 2c)
@@
The implications of the uderdetrimed fixed point solutions are surprising. Naively, one would expect that if the bottleneck is not too narrow, the projections of the error would maintain the necessary structure to guide the learning. Specifically, the Johnson–Lindenstrauss lemma  [17] suggests that as long as $r>d\log m$, the pairwise correlation structure of the error signal would be maintained. Nevertheless, our analysis shows that the solutions weights are not guaranteed to converge to the correct solution. Thus, in the case of low-rank feedback, Feedback Alignment (FA) is likely to fail. Furthermore, FA fails when $k<m$ even if the rank of $B$ is not restricted. The solution to this predicament, as we show next, is aligning the feedback weights with the data, ensuring that the network learns the correct representations.

### v4

# Efficiently Training Large Neural Networks With Low-Dimensional Error Feedback

@@ Abstract
Training large neural networks relies on propagating detailed error signals backward through every layer, a procedure that is computationally expensive and difficult to reconcile with biological teaching pathways. In the brain, instructive signals are typically sparse, indirect, and compressed, raising the question of whether a low-dimensional teaching signal can suffice for effective credit assignment without transmitting a full layerwise gradient. Building on Feedback Alignment, we introduce a low-dimensional feedback framework that decouples the backward and forward passes and explicitly controls the dimensionality of the error signal delivered to each layer while preserving high-dimensional feedforward representations. We develop a theory for the nonlinear learning dynamics induced by gradient descent in linear networks. Fixed, task-agnostic low-dimensional feedback can misassign credit and lead to spurious solutions, whereas allowing the feedback subspace to adapt toward task-relevant error directions recovers backpropagation-like learning dynamics. Guided by these predictions, we train convolutional networks and vision transformers to backpropagation-level performance with modest feedback dimensionality that scales with task dimensionality rather than network width, while reducing total training compute by over 20% without loss of accuracy. We further derive synapse-local rules that learn the relevant error subspace from pre- and postsynaptic signals, providing a biologically grounded mechanism for compressed teaching. Finally, feedback dimensionality reshapes receptive fields, revealing the teaching pathway as an inductive bias in representation learning.
@@
Training modern neural networks often depends on sending detailed error information backward through every layer, a process that is costly in time, energy, and memory. In contrast, the brain appears to rely on simpler teaching signals. We show that accurate learning can be driven by a compact feedback signal whose size is set by task demands rather than by the number of neurons or synapses. Building on this principle, we develop a training framework that matches the accuracy of the standard training method while being more efficient. We also describe local learning rules for constrained teaching pathways and show that these constraints shape features learned in early layers, offering a new route for modeling learning in brain circuits.
@@ 1.  Introduction
Training large neural networks depends on assigning credit across many layers and millions of parameters. The standard approach, gradient descent with error backpropagation, computes reverse-mode derivatives by transporting high-dimensional error information through every layer. This backward pass is not only a conceptual constraint but also a practical bottleneck because it requires manipulating and storing large intermediate signals whose size scales with hidden-layer width, leading to substantial memory, bandwidth, and compute costs during training [5, 11, 55]. While large forward representations are needed for expressivity, it is not obvious that the feedback pathway must also scale with model size.
This raises a basic efficiency question. Many tasks have low-dimensional outputs, so the loss gradient at the output layer is naturally low-dimensional, with an effective dimensionality tied to task complexity rather than to hidden-layer width. The open issue is whether deep networks can be trained by transmitting only a task-relevant, low-dimensional teaching signal to each layer, thereby reducing the overhead of the backward pass without sacrificing the benefits of high-dimensional feedforward representations.
This efficiency-motivated perspective is also consistent with biological circuitry. Biological circuits do not exhibit the synapse-wise, bidirectional connectivity and precise weight symmetry required by gradient backpropagation. Instead, experimentally observed instructive channels are typically indirect, not originating from the direct downstream neural population, and are often sparse and long-range, which implies a restricted teaching signal [42]. Notable examples include midbrain dopaminergic neurons [62, 45, 6], the olivocerebellar system [28, 47], cholinergic neuromodulation [33, 21, 36], and the noradrenergic system in the Locus Coeruleus [2, 7, 59, 67, 10]. Combined with empirical evidence that task-relevant population activity is often confined to low-dimensional manifolds [12, 16, 57, 22], these observations suggest that the brain’s teaching pathway is more plausibly indirect and compressed (low-dimensional) than a neuron-by-neuron gradient trace.
A parallel question arises in artificial learning. Existing biologically motivated alternatives to backpropagation, such as Feedback Alignment (FA) [38], decouple feedforward and feedback weights but still propagate full-dimensional error signals and often fall short of backpropagation in deep or complex architectures [4]. As a result, the role of error dimensionality as an explicit constraint on deep credit assignment has been difficult to isolate.
Here we develop a theoretical and algorithmic framework for training large networks with constrained, low-dimensional error feedback. We treat the error pathway as a controllable resource and ask which error subspace must be transmitted for effective learning, and how such a subspace can be learned with local plasticity rules. Our approach imposes a rank-$r$ constraint on feedback and learns a task-relevant low-dimensional teaching channel, thereby decoupling the backward pathway from the feedforward computation while preserving high-dimensional forward representations. We refer to this framework as low-dimensional feedback alignment (LDFA).
Our work makes four contributions:
  • We show that accurate credit assignment does not require full-dimensional backward signals. In a solvable linear theory, fixed low-rank feedback makes the stationarity conditions underdetermined and introduces spurious fixed points, and we characterize how the rank required for backpropagation-like convergence depends on task dimensionality.
  • We introduce a factorized low-rank feedback pathway together with learning rules that make the dimensionality of the layerwise teaching signal an explicit, tunable resource.
  • We demonstrate in deep nonlinear networks, including convolutional architectures and vision transformers, that near-backpropagation accuracy is achievable with ranks on the order of task complexity and largely independent of network width, and we quantify the resulting reduction in backward-pass compute.
  • We show that error dimensionality systematically shapes learned representations, including receptive-field structure in a simplified ventral-stream model.
Together, these results challenge the assumption that scalable credit assignment must involve transporting a full high-dimensional gradient through depth. Instead, they support a view in which learning can be driven by a learned low-dimensional teaching channel that preserves task-relevant error directions, yielding a computationally efficient training regime and a concrete hypothesis for how anatomically restricted feedback pathways could still support deep learning.
@@ 2.  Background and related work
[…]
Backpropagation (BP) [56] is the standard approach for training neural networks. It propagates the error backward through the network using $\delta_{l}=W_{l+1}^{\top}\delta_{l+1}\odot f^{\prime}(W_{l}\boldsymbol{h}_{l-1})$, and updates the weights using $\Delta W_{l}=-\eta\delta_{l}\boldsymbol{h}_{l-1}^{\top}$, where $\eta$ is the learning rate. However, this method requires an exact transposition of the forward weights, $W_{l+1}^{\top}$, which is biologically implausible [24, 15]. Moreover, BP tightly couples the error propagation with the forward pass, limiting the ability to explore how different properties of the error signal affect learning dynamics.
Feedback Alignment (FA) [38] was proposed to address the biological limitations of BP by replacing $W_{l+1}^{\top}$ with a fixed random matrix $B_{l}$. The error is computed as:
@@
$\Delta B_{l}=-\eta\boldsymbol{h}_{l-1}\delta_{l}^{\top}-\lambda B_{l},\qquad\Delta W_{l}=-\eta\delta_{l}\boldsymbol{h}_{l-1}^{\top}-\lambda W_{l},$ (2)
where $\lambda$ is a regularization parameter. Although this adaptive approach improves performance by better aligning forward and backward weights, it still requires high-dimensional error signals and and is less efficient than backpropagation. Furthermore, in section˜3, we show that this approach fails when the matrix $B$ is low-rank and the dimensionality of the error is constrained.
@@
In this work, we ask how limiting the dimensionality of the propagated error constrains credit assignment in large neural networks. We treat the error pathway as a controllable resource and study how its dimensionality shapes learning dynamics, computational cost, and the representations that emerge during training. To do so, we introduce a learning scheme in which each layer receives error through a low-rank feedback map, enabling explicit control over the dimensionality of the layerwise teaching signal (fig.˜1).
This focus differs from prior observations that learning dynamics or weight updates often exhibit low-dimensional structure: those results describe emergent compression but do not impose a bottleneck on the communicated error itself. The remainder of the paper develops a solvable linear theory of low-rank error feedback (section˜3), tests the predicted scaling in deep vision models (section˜4), derives synapse-local rules for learning the feedback pathway (section˜5), and then analyzes efficiency and representational consequences before closing with broader implications in the Discussion.
Figure 1: Illustration of different approaches for propagating error to hidden layers. From left to right: Backpropagation (BP) propagates error using the exact transposes of the forward weights. Feedback alignment (FA) replaces $W_{l}^{\top}$ with fixed random feedback matrices and relies on the forward weights aligning with this backward pathway during training. Low-dimensional feedback alignment (LDFA) constrains each feedback map to be low rank (e.g., $B_{l}=Q_{l}P_{l}$ with rank $r$), so each layer receives an $r$-dimensional teaching signal whose subspace is learned. Direct LDFA (dLDFA) extends LDFA by routing low-dimensional error projections to non-adjacent layers, allowing early layers to receive teaching signals directly from the output (or other higher) layers.
@@ 3.  Low-dimensional error feedback in linear networks
We consider a simple linear transformation problem with a low-dimensional structure, $\boldsymbol{y}=A\boldsymbol{x}$. Here, $\boldsymbol{x}\in\mathbb{R}^{n}$ represents the $n$-dimensional input, and $\boldsymbol{y}\in\mathbb{R}^{m}$ represents the target. The matrix $A$ is a rank-$d$ matrix defined as $A=\sum_{j=1}^{d}\boldsymbol{u}_{j}\boldsymbol{v}_{j}^{\top}$, where $\boldsymbol{u}_{j}\in\mathbb{R}^{n}$ and $\boldsymbol{v}_{j}\in\mathbb{R}^{m}$ are random Gaussian vectors, and we assume $d\ll n$. Our data set consists of $p$ training samples $\{\boldsymbol{x}^{\mu},\boldsymbol{y}^{\mu}\}_{\mu=1}^{p}$, with each input vector $\boldsymbol{x}^{\mu}$ being i.i.d. according to the standard normal distribution $\boldsymbol{x}^{\mu}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{1})$. The labels are given by $\boldsymbol{y}^{\mu}=A\boldsymbol{x}^{\mu}+\boldsymbol{\xi}^{\mu}$, where $\boldsymbol{\xi}^{\mu}$ is additive Gaussian noise with zero mean and unit variance.
@@
We apply Feedback Alignment (FA) to update $W_{1}$, which does not have direct access to the loss gradient. Instead of backpropagating the error through $W_{2}^{\top}$, we use a fixed low-rank feedback matrix $B$. This provides an alternative pathway for propagating the error signal to $W_{1}$.
@@
$\Delta W_{1}^{\mu}$ $=\eta B(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu})\boldsymbol{x}^{\mu\top},\text{ and }$ (4)
$\Delta W_{2}^{\mu}$ $=\eta(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu})\boldsymbol{x}^{\mu\top}W_{1}^{\top}.$
@@ 3.1 Learning dynamics
Our analysis extends the framework established by Saxe et. al. [60] to incorporate indirect feedback with constrained dimensionality. We begin by characterizing the task across the $p$ data points through the input-output covariance matrix, $\Sigma_{io}=\frac{1}{p}\sum_{\mu=1}^{p}\boldsymbol{y}^{\mu}(\boldsymbol{x}^{\mu})^{\top}$, which captures the correlation between input vectors $\boldsymbol{x}$ and output vectors $\boldsymbol{y}$. Performing Singular Value Decomposition (SVD), we obtain $\Sigma_{io}=USV^{\top}$, where $U\in\mathbb{R}^{m\times m}$ and $V\in\mathbb{R}^{n\times n}$ contain the left and right singular vectors, respectively, and $S\in\mathbb{R}^{m\times n}$ is a rectangular diagonal matrix of singular values. For sufficiently large $p$, the first $d$ singular values in $S$ correspond to the prominent directions in the data (i.e., the singular values of $A$), while the remaining singular values are $O(1/\sqrt{p})$ and dominated by noise.
@@
$\tau\frac{d\bar{W}_{1}}{dt}=\bar{B}(S-\bar{W}_{2}\bar{W}_{1}),\qquad\tau\frac{d\bar{W}_{2}}{dt}=(S-\bar{W}_{2}\bar{W}_{1})\bar{W}_{1}^{\top},$ (5)
@@
where $\eta=dt/\tau$. This continuous form captures the dynamics of the learning process, allowing us to study it from a dynamical systems perspective. By analyzing these equations, we can identify fixed points and evaluate their stability, providing insight into how the network converges and learns under constrained feedback.
@@
Figure 2: Learning dynamics and component alignment in linear networks. (a) Schematic of the network architecture with input dimension $n=128$, hidden layer size $k=64$, and output dimension $m=64$. The feedback matrix $B$ is factorized as $B=QP$ and constrained to rank $r$. (b, c) Comparison of theoretical predictions (dashed) and numerical simulations (solid) for low-rank Feedback Alignment (FA), updating $Q$ with $P$ fixed [1], with $r=m=64$ and $r=8<m$, respectively. The $y$-axis reports the normalized mode overlap $\Lambda_{i}=u_{i}^{\top}W_{2}W_{1}v_{i}/S_{ii}$ for each singular component ii. (d, e) Same as (b, c) but training both $Q$ and $P$ using Eqs. (9). (f) With $P$ fixed, overlaps increase on average (bold), but the leading singular components do not reliably reach $\Lambda_{i}\approx 1$ at low rank. (g) Training $P$ aligns the feedback subspace with the evolving error, yielding near-complete recovery of the top-$r$ components and improved convergence.
Stationary solutions for training.
@@
Training halts when the right-hand side of the weight dynamics in (5) vanishes, that is, when the system reaches a stationary point. At such a point, the update equation for $\bar{W}_{1}$ yields
$\bar{B}\,(S-\bar{W}_{2}\bar{W}_{1})=0\iff S_{jj}\,\boldsymbol{B}_{:,j}=\sum_{i=1}^{m}\boldsymbol{B}_{:,i}\,(\bar{W}_{2}\bar{W}_{1})_{ij}\qquad\forall j,$ (6)
where $\boldsymbol{B}_{:,j}\in\mathbb{R}^{k}$ is the $j$-th column of $\bar{B}$ and $S_{jj}$ is the $j$-th singular value of $\Sigma_{io}$. Equation (6) constrains the mismatch between the target map $S$ and the learned input–output map $\bar{W}_{2}\bar{W}_{1}$ only through the subspace spanned by the columns of $\bar{B}$.
A convenient way to quantify how well the learned map captures each singular mode of the data is to project $\bar{W}_{2}\bar{W}_{1}$ into the singular-vector basis of $\Sigma_{io}=USV^{\top}$. We therefore define the (diagonal) mode overlap
$\Lambda_{i}=\frac{u_{i}^{\top}(\bar{W}_{2}\bar{W}_{1})v_{i}}{S_{ii}}=\frac{(U^{\top}\bar{W}_{2}\bar{W}_{1}V)_{ii}}{S_{ii}},$ (7)
where $u_{i}$ and $v_{i}$ are the $i$-th columns of $U$ and $V$. The overlap $\Lambda_{i}$ is a normalized measure of how strongly the learned input–output map expresses singular mode $i$ of $\Sigma_{io}$, with $\Lambda_{i}=1$ corresponding to perfect recovery of that mode.
When $\mathrm{rank}(\bar{B})=r=m$, the feedback pathway constrains all $m$ directions, and the stationary conditions select the backpropagation solution $\bar{W}_{2}\bar{W}_{1}=S$, as in the exact linear-network dynamics of Saxe et al. [60] (Fig. 2b).
Crucially, when $r<m$, the constraint (6) is generically underdetermined because $\bar{B}$ can span at most $r$ independent directions. As a result, there can be infinitely many stationary solutions, and $\bar{W}_{2}\bar{W}_{1}$ need not match the full singular structure of $\Sigma_{io}$, so that some mode overlaps $\Lambda_{i}$ can deviate substantially from $1$ (Fig. 2c).
@@
The implications of underdetermined fixed-point solutions are surprising. Naively, one would expect that if the bottleneck in the feedback is not too narrow, the projections of the error would maintain the necessary structure to guide the learning. Specifically, the Johnson–Lindenstrauss lemma  [31] suggests that as long as $r>d\log m$, the pairwise correlation structure of the error signal is maintained. Nevertheless, our analysis shows that the solution weights are not guaranteed to converge to the correct solution. Thus, in the case of low-rank feedback, Feedback Alignment (FA) is likely to fail. Furthermore, FA fails when there is a bottleneck, i.e., $k<m,n$, even if the rank of $B$ is not restricted. The solution to this predicament, as we show next, is aligning the feedback weights with the data, ensuring that the network learns the correct representations.